OpenStudy (anonymous):

Auditions for a new, upcoming band in Chicago were held with 102 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to -6, how many women auditioned?

OpenStudy (anonymous):

m+w=102 2m-3w=-6 2(102-w)-3w=-6 -5w=-210 w=42 m=60 -

OpenStudy (stamp):

Ok well there is your answer, I hope you learned something.